OpenStudy (butterflydreamer):

Newton's Method Q. Taking x = 1.5 as a first approximation, use one step of Newton's method to find a better approximation to this root, ln x = 2 - x

OpenStudy (phi):

\[ x_{n+1}= x_n - \frac{f(x_n)}{f'(x_n) } \]

OpenStudy (butterflydreamer):

so would that mean x1 = 1.5 and we're trying to find x2 ? But how do we get f(x) ? Would it just be f(x) = ln x - 2 + x ?

OpenStudy (phi):

yes

OpenStudy (phi):

you need a calculator to evaluate f(1.5) f'(x) = 1/x + 1 and f'(1.5) = 5/3 the new estimate for x will be = 1.5 - 3/5 * ( ln(1.5) + 1.5 -2)

OpenStudy (butterflydreamer):

thankk you :D @phi . I ended up with x2 = 1. 56

OpenStudy (dan815):

do u know how newtons method works

OpenStudy (dan815):

okay suppose you are given some function of x, we first rewrite in the form f(x) =0 okay lets just suppose this f(x) is a parabola for simplifcity sake, u will be able to see why it extend all almost all other polynomial functions

OpenStudy (dan815):